# JumpStarter: Getting Started on Personal Goals with AI-Powered Context Curation

### v2

@@ Abstract.
Everyone aspires to achieve personal goals. However, getting started is often complex and daunting, especially for large projects. AI has the potential to create plans and help jumpstart progress, but it often lacks sufficient personal context to be useful. We introduce JumpStarter, a system that uses AI-powered context curation to create action plans and draft personalized working solutions. JumpStarter assists users by posing questions to elicit relevant context, breaking down goals into manageable steps, and selecting appropriate context to draft working solutions for each step. A technical evaluation indicates that context curation results in plans and working solutions of higher quality. A user study demonstrates that compared to ChatGPT, JumpStarter significantly reduces users’ mental load and increases their efficiency in kickstarting personal projects. We discuss the design implications of AI-powered context curation to facilitate the use of generative AI in complex problem-solving.
@@
Figure 1. JumpStarter helps users get started on their personal goals through AI-powered context curation. It first takes the user’s goal and gathers context for the goal. It then breaks down the goal into actionable subtasks. For each subtask, it helps users choose relevant context and draft working solutions. It also aids users in refining these drafts by eliciting further context.
@@ 1. Introduction
Achieving personal goals is a common aspiration. Personal projects are part of everyday life, ranging from moving to a new apartment, to starting a YouTube channel, to applying for PhD programs. Getting started on these projects is crucial for moving from mere intention to actual fulfillment (Sheeran and Webb, 2016; Gollwitzer, 1996). However, getting started on complex goals can be daunting and challenging. When faced with a new project, one needs to engage in sensemaking, construct a mental model of the process, and adapt it to their specific situation to figure out the actionable steps, all of which can be cognitively demanding (Johnson-Laird, 1983). Another major challenge is initiating concrete actions. For example, there is a significant gap between knowing one needs to write a request email and actually being ready to write it (Sheeran and Webb, 2016). This shift requires concretization—formulating a specific request or language and gathering the necessary information. Identifying and utilizing the right resources to operationalize the plan can be a significant hurdle to getting started.
@@
Generative AI has the potential to create plans and help jumpstart progress, but it often lacks the personal context needed to be useful. Everyone has unique situations and requirements that need to be considered when embarking on personal projects. For example, to get started on the goal of “applying for a PhD,” it is important to consider preferred research areas and personal background. Moreover, a complex goal like this must be further decomposed to become actionable. It involves juggling multiple tasks, such as researching universities and requesting recommendation letters, each of which requires further strategic planning and coordination with other tasks. For example, requesting recommendation letters is itself a goal that needs further breakdown—one must first identify recommenders, gather supporting documents, send them individualized email drafts, and follow up accordingly. Popular AI tools like ChatGPT allow users to provide context throughout their interactions. However, it can be mentally demanding for users to recall and articulate the necessary context relevant to the task and to frame the prompt accordingly. Ideally, ChatGPT would remember all the context provided by users in the chat and select the most appropriate ones to support each task. However, this capability is limited by the inherent restrictions on context window size (Chen et al., 2023; Peng et al., 2023; Yen et al., 2024), and the process is not transparent to users. Moreover, linear chat interfaces like ChatGPT lack the structure to help people systematically plan their roadmap toward their goals. Although people can ask ChatGPT to break down the steps further, it is easy to lose track of their progress and the context of each task.
@@
In the psychology literature, the management of personal goals is often linked to executive functions such as planning, task initiation, and working memory (Diamond, 2013). Planning involves a step-by-step breakdown of a broader goal (Kokkalis et al., 2013). This can be challenging as it requires identifying key tasks, estimating realistic timelines, and foreseeing potential obstacles, all demanding significant effort (Teevan et al., 2016). While having a plan is crucial, it can sometimes remain abstract and lack the specificity and granularity required for tangible actions (Gollwitzer, 1996; Kokkalis et al., 2013). Initiating concrete tasks demands heightened focus and meticulous attention to detail, which significantly increases cognitive load (Sweller, 2011). Effective execution of such tasks necessitates maintaining an active working memory, which aids in retaining and manipulating the relevant information needed to manage the task (Baddeley, 1992). Additionally, the Ovsiankina effect (Rickers-Ovsiankina, 1928) indicates that people are more likely to finish the task once they initiate it, underscoring the critical role of action initiation in goal completion.
@@
To better support people in getting started on their personal goals, we introduce a system called JumpStarter (see Figure 1). Through AI-powered context curation, the system assists users by 1) decomposing tasks into hierarchical steps and 2) drafting working solutions. Here, the “working solution” refers to any tangible result that enables users to take real-world action. For example, a working solution could be a draft email to send to a recommender or a study schedule for the GRE. The process begins when a user inputs their personal goal. The system then poses questions to elicit the relevant context for the goal. Based on user responses, the system breaks down the goal into subtasks. The user can then start exploring each subtask. The system notifies users if it detects that the subtasks are not actionable enough, prompting further breakdown. If the user opts to decompose further, the system assists in generating subtasks. It also helps detect if “forking” is necessary, which involves breaking down the task based on different entities (such as sending requests to different recommenders), focusing on one at a time. If the user is ready to start on a task, the system selects the relevant context and generates working solutions that incorporate this context. If the user is not satisfied with the suggested working solution, they can request more elicitation questions or provide their own prompts to refine it. Users can select parts of the suggested solution and attach them to the subtask, and the system will add them to the context pool. The user can then move on to the next subtask to explore.
@@
  • JumpStarter, an AI-powered system for helping people get started on their personal goals with three main automated features: 1) context curation (including context elicitation, context saving, and context selection), 2) task decomposition and 3) working solution drafting.
@@
  • A user study showing that compared to ChatGPT, JumpStarter significantly reduces users’ mental load and increases their efficiency in kickstarting personal projects.
We conclude by discussing the design implications of AI-powered context curation to facilitate the use of generative AI in complex problem-solving.
@@ 2.2. Solutions to Support Action Planning
A popular online tool called Goblin Tools MagicTodo [^1] https://goblin.tools/ uses AI techniques to support people in breaking down their tasks layer by layer, but the accuracy seems limited. A recent HCI work, ExploreLLM (Ma et al., 2024), also explores helping users decompose complex tasks into subtasks with LLMs. Their study shows that a schema-like structure is useful for people to do planning tasks, with ChatGPT as a baseline. The system also provides a dedicated input box to prompt users to enter their personal context and preferences for personalization. However, ExploreLLM only examines one layer of task decomposition and relies on users to provide context themselves. In JumpStarter, we explore multi-layer task decomposition and intelligent context curation throughout the process.
@@ 2.3. Solutions to Support Action with Context Curation
Curating appropriate context relevant to a task is essential for enabling people to take action on it. However, people often need to juggle multiple tasks in daily life, and context switching is never effortless (Kersten and Murphy, 2006), hindering people’s ability to make progress. Therefore, task-specific context management is important and beneficial to support.
Previous system works focus on help people do task-centric information management. For example, UMEA (Kaptelinin, 2003) helps users create and manage project-specific work context by providing users dedicated project spaces and monitoring user activities in a desktop environment. Mylar (Kersten and Murphy, 2006) introduces a task context model to help programmers find and identify the information relevant to their task at hand, which is built using the task interaction history. Personal Project Planner (Jones et al., 2008) introduces an in-context create feature that enables users to produce documents and emails within a project plan, automatically inserting hyperlinks as context to the task. Similarly, Towel (Conley and Carpenter, 2007) allows users to drag any semantic object (project, email, person etc.) onto the to-do list, saving them as resources under the to-do item.
To better support users in taking action on their tasks, the existing context is often not sufficient, and the system may need to elicit more context from the user. For example, PExA (Myers et al., 2007) introduces an intelligent personal assistant that helps people manage their tasks by suggesting context-appropriate follow-up questions. The use of large language models has also been explored to support context elicitation. For example, the OPEN framework (Handa et al., 2024) uses Bayesian Optimal Experimental Design to choose informative queries and employs language models to formulate these queries in natural language, enhancing preference elicitation accuracy. Generative Active Task Elicitation (GATE) (Li et al., 2023) leverages language models to interactively generate questions and scenarios, guiding users to clarify and refine their intended outcomes in tasks, which helps elicit more precise and contextually rich information from users. In JumpStarter, we explore generative AI-powered context curation, which covers task-specific context elicitation, context saving, and context selection, in the scenario of getting people started on complex personal projects.
@@ 3. JumpStarter System
JumpStarter is an interactive system that helps users get started on their personal projects with AI-powered context curation. The system takes the user-specified goal as the input, and outputs plans and working solutions for the user. Here, a working solution refers to any tangible result that can assist users in taking real-world actions related to a task. For instance, a working solution for requesting recommendation letters might be an email draft to solicit letters from advisors. Similarly, for passing the GRE test, a working solution could include a study schedule. JumpStarter is powered by GPT-4222We used gpt-4-turbo throughout the system. and implemented with a Flask/Python web framework. In this section, we describe how our system works with an example walk-through, and the implementation details of the system.
@@ 3.1. System Description
Context Elicitation for Goal
Based on John’s goal, the system uses the LLM to generate questions to elicit context from him (see Figure 3). For John, the system generates three questions: ”Can you provide your CV or Resume?”, ”Which universities or programs are you considering for your PhD in NLP?”, and ”Do you have letters of recommendation, or do you need guidance on how to obtain them?”. Since John does not know which universities he might be applying to, and he has not obtained the recommendation letters, he only uploads his CV to the first question. Then, he clicks the Let’s get started button to proceed. This first step aims to clarify John’s goal by eliciting the necessary context that will be used globally throughout John’s interaction with JumpStarter, ensuring that the plans and working solutions are created based on John’s situations.
Figure 3. JumpStarter generates questions to elicit necessary context from users to clarify the goal. The user uploads his CV and adds his own context Preferred location: United States.
Subtask Suggestion
@@
Subtask Detection
After reviewing the initial subtask tree, John thinks the list generally makes sense and proceeds to click on the first node: Identify Potential PhD Programs: Research and create a list of universities and faculty members who specialize in NLP. As shown in Figure 5, the node is highlighted, and its title and description are shown on the top right of the interface. At the same time, the system shows a message as it detects this task needs to be further decomposed to make it actionable. The system offers two options for the user to select: the recommended option Decompose the task and the alternative option Directly start drafting. Clicking on the first button will further break down the selected task into several subtasks, while clicking on the second will generate a working solution for the task. Following the recommendation, John clicks the button Decompose the task, after which the subtasks for Identify Potential PhD Programs are generated, as shown in Figure 5.
@@
Working Solution Drafting
John reviews all subtasks of Identify Potential Programs and proceeds to click on the first subtask Research Universities and Programs, which is about “Identify universities that have strong PhD programs in NLP. Use academic rankings, university websites, and research publications to compile a list of potential programs.” JumpStarter deems this subtask actionable and suggests John directly start drafting the working solution for it. Following the suggestion, he obtained the initial working solution that contains a list of universities offering PhD programs in NLP. Upon reviewing the list, John realizes it includes many global institutions. However, he prefers schools in the midwest of the United States. To iterate, he enters ”I want schools in midwest of US” (see part D in Figure 2). After reading it, John is happy to have this list as the first draft of his school application list. Note that John is provided with three alternatives to further improve the draft, as shown in part D of Figure 2: 1) regenerate; 2) add more context and regenerate; and 3) specify details and iterate on the current draft. Subsequently, John copies and pastes the draft into the working solution section of the task Research Universities and Programs (see part B in Figure 2) and saves it, which appears as a file icon on the tree (see part A in Figure 2) with the subtask crossed out.
@@
Forking Detection and Context Selection
After obtaining the initial working solution for the first subtask, John moves on to the next subtask titled Identify Faculty Members. JumpStarter suggests that the subtask lacks the specificity needed for direct action and suggests further breakdown. John agrees and clicks the Decompose the task button. Following this, JumpStarter determines that the subtask should be forked to address the specifics of each program identified earlier. It then selects and presents the relevant context, specifically the working solution from the previously completed step Research Universities and Programs that includes a list of universities (see Figure 6(a)). John agrees with the system’s suggestion and clicks the Accept button to proceed with forking using the selected context. Note that if John disagrees with the suggested context, he can select other context variables from a dropdown button; alternatively, he can also Decline and proceed with breakdown if he thinks that the subtask should not be forked. The updated subtask tree after forking is displayed in Figure 6(b). With the new subtasks of evaluating different programs, John proceeds to explore each program in greater detail by clicking on them sequentially.

### v3

# JumpStarter: Getting Started on Personal Goals with Adaptive Personal Context Curation

@@ Abstract.
Everyone aspires to achieve personal goals. However, getting started is often complex and daunting, especially for large projects. AI has the potential to create plans and help jumpstart progress, but it often lacks sufficient personal context to be useful. Personal context includes background information, preferences, and resources that can help tailor AI responses to individuals and their goals. We introduce JumpStarter, an AI-powered system that creates action plans and drafts by adaptively soliciting and integrating personal context. JumpStarter assists users by posing questions to elicit relevant context, breaking down goals into manageable steps based on context, and identifying relevant context to draft working solutions. A technical evaluation indicates that context curation results in plans and working solutions of higher quality. A user study demonstrates that compared to ChatGPT, JumpStarter significantly reduces mental load and increases efficiency for users. We discuss the design implications of adaptive personal context curation to better support users in complex problem solving.
@@
Figure 1. JumpStarter helps users get started on their personal goals through adaptive personal context curation. It first takes the user’s goal and elicits context for the goal. It then breaks down the goal into actionable subtasks. For each subtask, it helps users select relevant context and draft working solutions. It also aids users in refining these drafts by eliciting further context.
@@ 1. Introduction
Achieving personal goals is a common aspiration. Personal projects are part of everyday life, ranging from moving to a new apartment, to starting a YouTube channel, to applying for PhD programs. Getting started on these personal projects is crucial to turn intention into actual fulfillment (Sheeran and Webb, 2016; Gollwitzer, 1996). However, getting started on complex goals can be daunting and challenging. When faced with a new project, one needs to engage in sensemaking: constructing a mental model of the process and adapting it to their specific situation to figure out the actionable steps, all of which can be cognitively demanding (Johnson-Laird, 1983). Another major challenge is initiating concrete actions. For example, there is a significant gap between knowing when one needs to write a request email and actually being ready to write it (Sheeran and Webb, 2016). This shift requires concretization—such as defining the exact purpose of the request, deciding on the appropriate tone, and gathering the necessary information like relevant background details and supporting documents. Identifying and utilizing the right resources to operationalize the plan can be a significant hurdle to getting started.
@@
Generative AI has the potential to create plans and help jumpstart progress, but it often lacks the personal context needed to be useful. Everyone has unique situations and requirements to consider when embarking on personal projects. For example, to get started on the goal of “applying for a PhD,” it is important to consider personal context, such as one’s academic background and geographic preferences. Moreover, a complex goal like this must be further decomposed to become actionable. It involves multiple tasks, such as researching universities and requesting recommendation letters, each of which requires further planning and coordination with other tasks. For example, requesting recommendation letters is itself a sub-goal that needs further breakdown—one must first identify recommenders, gather supporting documents, draft individualized request emails, send the emails, and follow up accordingly. Each of these tasks demands adaptive curation of relevant context to craft personalized solutions. For example, when drafting emails to recommenders, it is essential to solicit relevant context to ensure that the drafts are personalized to the recipient, which might include one’s relationship with the recommender, the programs being applied to, and aspirations for doctoral studies.
Popular AI tools like ChatGPT allow users to provide context throughout their interactions. However, it can be mentally demanding for users to recall and articulate the necessary context relevant to the task and to structure all the information within the bounds of prompts and chat history. Ideally, ChatGPT would remember all the context in the chat and automatically integrate the most relevant ones to support each task. However, this capability is limited by the inherent restrictions on context window size (Chen et al., 2023; Peng et al., 2023; Yen et al., 2024), and the context integration process is not transparent to users. Moreover, linear chat interfaces like ChatGPT lack the structure to help people systematically plan their roadmap toward their goals. Although people can ask ChatGPT to break down the steps further, it is easy to lose track of their progress and the context of each task.
@@
Furthermore, theories about personal goals link their successful implementation to executive functions such as planning, task initiation, and working memory (Diamond, 2013), so approaches should be grounded in these functions. Planning involves a step-by-step breakdown of a broader goal (Kokkalis et al., 2013). This can be challenging as it requires identifying key tasks, estimating realistic timelines, and foreseeing potential obstacles, all of which demand significant effort (Teevan et al., 2016). While having a plan is crucial, it can sometimes remain abstract and lack the specificity and granularity required for tangible real-world actions (Gollwitzer, 1996; Kokkalis et al., 2013). Initiating concrete tasks demands heightened focus and meticulous attention to detail, which significantly increases cognitive load (Sweller, 2011). Effective execution of such tasks requires an active working memory, which aids in retaining and manipulating the relevant information needed to manage the task (Baddeley, 1992). Additionally, the Ovsiankina effect (Rickers-Ovsiankina, 1928) indicates that people are more likely to finish the task once they initiate it, underscoring the critical role of action initiation in goal completion.
@@
To better support people in getting started on their personal goals, we introduce a system called JumpStarter (see Figure 1). Through mechanisms that adaptively solicit and integrate personal context, the system assists users by 1) decomposing tasks into hierarchical steps and 2) drafting working solutions. Here, the “working solution” refers to any tangible result that enables users to take real-world action. For example, a working solution could be a draft email to send to a recommender or a study schedule for the GRE. The process begins when a user inputs their personal goal. The system then poses questions to elicit the relevant context for the goal. Based on the user’s responses, the system breaks down the goal into subtasks the user can explore. If the system detects that a subtask is not actionable enough, it prompts the user, who can break it down further into additional subtasks. When the user is ready to work on a task, the system selects the relevant context and generates working solutions. If the user is not satisfied with the suggested solution, they can request more elicitation questions or provide their own prompts for refinement. Users can modify the suggested solution, and also save it as context to be used by the system in further subtasks. The user can then move to explore the next subtask, moving towards the completion of their goal.
@@
  • JumpStarter, an AI-powered system for helping people get started on their personal goals with three main automated features: 1) adaptive personal context curation (including eliciting, saving, and selecting context), 2) hierarchical task decomposition, and 3) working solution generation.
@@
  • A user study showing that compared to ChatGPT, JumpStarter significantly reduces mental load and increases efficiency in getting started on personal projects.
We conclude by discussing the design implications of adaptive personal context curation to guide the design of future systems that support complex problem solving.
@@ 2.2. Solutions to Support Action Planning
A popular online tool called Goblin Tools MagicTodo [^1] https://goblin.tools/ uses AI techniques to support people in breaking down their tasks layer by layer. Throughout our experiments with MagicTodo, we identified significant limitations, particularly in determining the appropriate level of task breakdown granularity. In theory, tasks can be endlessly decomposed into trivial steps, such as opening a browser. The challenge is to find a manageable granularity that remains meaningful for users. To address this, JumpStarter examines subtask detection to help users decide whether a task requires further breakdown while still allowing for personal control. A recent HCI work, ExploreLLM (Ma et al., 2024), also explores helping users decompose complex tasks into subtasks with LLMs. Their study shows that a schema-like structure is useful for people to do planning tasks, with ChatGPT as a baseline. The system also provides a dedicated input box to prompt users to enter their personal context and preferences for personalization. However, ExploreLLM only examines one layer of task decomposition and relies on users to provide context themselves. In JumpStarter, we explore multi-layer task decomposition and adaptive personal context curation throughout the process.
@@ 2.3. Solutions to Support Action with Context Curation
Context curation is a fundamental design concept that impacts productivity. It involves eliciting, saving, and selecting relevant information to tailor solutions for specific tasks and individuals (Kerne et al., 2014). The idea of ”readiness-to-hand” (Heidegger, 2010) illustrates how we perceive tools as extensions of our abilities, emphasizing the close intertwining of cognition and surroundings. For example, mise-en-place (Weisberg et al., 2014), a strategic arrangement of tools and ingredients before cooking, demonstrates how well-curated contexts can enhance efficiency. According to cognitive theories, our thoughts and actions are adaptive responses to our environment, underscoring the deep connection between human cognition and context (Winograd, 1986). This structural coupling highlights the significance of context curation in improving productivity and interactions with technology.
Context elicitation has been studied in the NLP field. Research in proactive information-seeking dialogue (Deng et al., 2023; Zhang et al., 2023) focuses on approaches to asking clarifying questions (Zhang et al., 2024; Malaviya et al., 2024) and eliciting user preferences (Zhang et al., 2018; Handa et al., 2024; Li et al., 2023) to better meet users’ information needs. However, this body of research often centers on pure information-seeking tasks, such as web searches or online purchases, and the experiments are often conducted in simulated chat environments. JumpStarter broadens this scope by facilitating interactions with real users across a wider variety of real-world tasks in a structured manner. Moreover, it advances this research by eliciting context beyond typical question-and-answer exchanges and incorporating personal files into the process, making context elicitation more practical.
Context saving and selection were studied in task-centric information management systems. In terms of context saving, UMEA (Kaptelinin, 2003) helps users create and manage project-specific work context by providing dedicated project spaces and monitoring user activities within a desktop environment. The Personal Project Planner (Jones et al., 2008) introduces an in-context creation feature that enables users to produce documents and emails within a project plan, automatically inserting hyperlinks to provide context to the task. Similarly, Towel (Conley and Carpenter, 2007) allows users to drag any semantic object (e.g., project, email, person) onto the to-do list, saving them as resources under the to-do item. Context selection has been particularly examined in programming scenarios. For example, Mylar (Kersten and Murphy, 2006) introduces a task context model to help programmers find and identify information relevant to their current task by utilizing task interaction history. JumpStarter not only streamlines the integration of context into users’ tasks but also dynamically optimizes task execution, transforming static context management into a proactive and adaptive workflow catalyst.
@@ 3. JumpStarter System
JumpStarter (see Figure 2) is an interactive system that helps users get started on their personal projects with adaptive personal context curation. The system takes the user-specified goal as the input, and outputs personalized plans and working solutions for the user. Here, a working solution refers to any tangible result that can assist users in taking real-world actions related to a task. For instance, a working solution for requesting recommendation letters might be an email draft to send to the advisor. Similarly, for passing the GRE test, a working solution could include a study schedule.
Overall, JumpStarter utilizes a hierarchical tree structure. We chose this approach to organize tasks because it offers a clear, logical flow that naturally mirrors the breakdown of complex goals into manageable components. This structure facilitates easy tracking of task dependencies and progression. Features such as collapsing or expanding nodes allow for customizable views, ranging from simple outlines to detailed oversight. Additionally, the ability to attach documents directly to task nodes enhances organization and accessibility, making this approach particularly effective for managing personal projects.
JumpStarter is powered by GPT-4222We used gpt-4-turbo throughout the system. and implemented with a Flask/Python web framework. In this section, we describe how our system works with an example walk-through, and the implementation details of the system.
@@ 3.1. System Description
3.1.1. Context Elicitation for Goal
Based on John’s goal, the system uses the LLM to generate questions to elicit context from him (see Figure 3). For John, the system generates three questions: ”Can you provide your CV or Resume?”, ”Which universities or programs are you considering for your PhD in NLP?”, and ”Do you have letters of recommendation, or do you need guidance on how to obtain them?” Since John does not know which universities he might be applying to, and he has not obtained the recommendation letters, he only uploads his CV to the first question. Then, he clicks the Let’s get started button to proceed. This first step aims to clarify John’s goal by eliciting the necessary context that will be used globally throughout John’s interaction with JumpStarter, ensuring that the plans and working solutions are created based on John’s situations.
Figure 3. JumpStarter generates questions to elicit context from users to clarify the goal. The user uploads his CV.
3.1.2. Subtask Suggestion
@@
3.1.3. Subtask Detection
After reviewing the initial subtask tree, John thinks the list generally makes sense and proceeds to click on the first node: Identify Potential PhD Programs. As shown in Figure 5, the node is highlighted, and its title and description are shown on the top right of the interface. At the same time, the system shows a message as it detects this task needs to be further decomposed to make it actionable. The system offers two options for the user to select: the recommended option Decompose the task and the alternative option Directly start drafting. Clicking on the first button will further break down the selected task into several subtasks, while clicking on the second will generate a working solution for the task. Following the recommendation, John clicks the button Decompose the task, after which the subtasks for Identify Potential PhD Programs are generated, as shown in Figure 5.
3.1.4. Working Solution Drafting
John reviews all subtasks of Identify Potential Programs and proceeds to click on the first subtask Research Universities and Programs. JumpStarter deems this subtask actionable and suggests John directly start drafting the working solution for it. Following the suggestion, he obtained the initial working solution that contains a list of universities offering PhD programs in NLP. Upon reviewing the list, John realizes it includes many global institutions. However, he prefers schools in the midwest of the United States. To iterate, he enters ”I want schools in midwest of US” (see Figure 2D). After reading it, John is happy to have this list as the first draft of his school application list. Note that John is provided with three alternatives to further improve the draft, as shown in Figure 2D: 1) regenerate; 2) add more context and regenerate; and 3) specify details and iterate on the current draft. Subsequently, John copies and pastes the draft into the working solution section of the task Research Universities and Programs (see Figure 2B) and saves it, which appears as a file icon on the tree (see Figure 2A) with the subtask crossed out.
@@ 3.1.4. Working Solution Drafting
3.1.5. Forking Detection and Context Selection
After obtaining the initial working solution for Research Universities and Programs, John moves on to the next subtask titled Identify Faculty Members. JumpStarter suggests that the subtask lacks the specificity needed for direct action and suggests further breakdown. John agrees and clicks the Decompose the task button. Following this, JumpStarter determines that the subtask should be forked to address the specifics of each program identified earlier. It then selects and presents the relevant context, specifically the working solution of the previously completed step Research Universities and Programs that includes a list of universities (see Figure 6(a)). John agrees with the suggestion and clicks the Accept button to proceed with forking using the selected context. Note that if John disagrees with the suggested context, he can select other context variables from a dropdown button; alternatively, he can also Decline and proceed with breakdown if he thinks that the subtask should not be forked. The updated subtask tree after forking is displayed in Figure 6(b). With the new subtasks of evaluating different programs, John proceeds to explore each program in greater detail by clicking on them sequentially.
3.1.6. Context Selection for Working Solution Generation
Next, John decides to explore the task Get Recommendation Letters and decomposes it as the system suggests. John selects the first subtask Compile a list of recommenders. JumpStarter uses his CV uploaded in the first step to generate a draft of recommenders: Prof. Blake White, Prof. Julian Deng, and Dr. Alice Feng. John likes the draft and saves it. Subsequently, John moves on to the subtask Reach out to potential recommenders, and JumpStarter breaks down the task based on the recommenders identified in the previous step. John then selects the subtask Reach Out to Potential Recommenders: Prof. Blake White, which requires “Send a personalized email to Prof. Blake White about your PhD aspirations.“ In addition to John’s CV, JumpStarter suggests two more relevant context: the information of each recommender (contains the relationship between Prof. Blake White and John) and the list of potential schools, as shown in Figure 7(a). Following the suggestion, John receives the initial email draft.
(a) Context selection for working solution generation.
(b) Context elicitation for working solution iteration.
Figure 7. Context selection and elicitation for drafting and iterating the working solution of the subtask Reach Out to Potential Recommenders: Prof. Blake White.
3.1.7. Context Elicitation for Working Solution Iteration
After reviewing the initial draft of the email for Prof. Blake White, John decides it needs improvement, though he is unsure how to proceed. He clicks the ”Add Context and Regenerate” button, prompting JumpStarter to generate several questions designed to draw more context from him (See Figure 7(b)). John then includes details about specific projects and papers he collaborated on with Prof. White and regenerates a refined draft. John likes the new version because it highlights specific aspects of his past collaborations with Prof. White, giving it a more personalized touch.